OpenStudy (anonymous):

Algebra 2 PLEASE HELP!!! I'm so confused... (X+Y)(X+Y) X^2 + XY + YX +Y^2 Supposedly (X^2 +2XY + Y^2) Should the 2XY be 2X2Y? Because there's two of each not just two of the "X".

OpenStudy (freckles):

(x+y)(x+y) is x^2+2xy+y^2 an example: say x=2 and y=3 (2+3)(2+3)=(5)(5)=25 right? well what if we multiply (2+3)(2+3) like you did above: \[(2+3)(2+3) \\ 2(2+3)+3(2+3) \\ 2^2+2(3)+3(2)+3^2 \\ 2^2+2 \cdot 2(3)+3^2 \\ 4+4(3)+9 \\ 4+12+9 \\ 16+9 \\ 25\]

rishavraj (rishavraj):

(a + b)(a + b) = (a + b)^2 = a^2 + b^2 + 2ab

OpenStudy (freckles):

like x+x=2x right? so I said 2(3)+2(3)=2*2(3)